OpenStudy (anonymous):

I will give you a medal if can help me Solve the system using the elimination method. 6x+12y=-7 x+2y=2 Please show work

OpenStudy (anonymous):

1(6x+12y=-7) 6(x+2y=2) 6x + 12y = -7 6x + 12y = 12 Notice if you subtract you get -7=12 which is false. There are NO SOLUTIONS, as they are parallel lines.